# for Improving Reasoning

### v1

@@ Abstract
The reasoning steps generated by LLMs might be incomplete, as they mimic logical leaps common in everyday communication found in their pre-training data: underlying rationales are frequently left implicit (unstated). To address this challenge, we introduce Rationalyst, a model for process-supervision of reasoning based on pre-training on a vast collection of rationale annotations extracted from unlabeled data. We extract 79k rationales from web-scale unlabelled dataset (the Pile) and a combination of reasoning datasets with minimal human intervention. This web-scale pre-training for reasoning allows Rationalyst to consistently generalize across diverse reasoning tasks, including mathematical, commonsense, scientific, and logical reasoning. Fine-tuned from LLaMa-3-8B, Rationalyst improves the accuracy of reasoning by an average of 3.9% on 7 representative reasoning benchmarks. It also demonstrates superior performance compared to significantly larger verifiers like GPT-4 and similarly sized models fine-tuned on matching training sets. 111 Our code, data, and model can be found at this repository: https://github.com/JHU-CLSP/Rationalyst
Rationalyst: Pre-training Process-Supervision
for Improving Reasoning
@@ 1 Introduction
Rationales play a crucial role in human reasoning and its accuracy (Rips, 1994; Mercier and Sperber, 2011). In reasoning problems, having accurate rationales often correlates with accurate outcomes (Tversky et al., 1982; Davis, 1984). This importance of rationales extends to Large Language Models (LLMs) as well. Wei et al. (2022) were among the first to show that generating chain-of-thought rationales significantly improves LLMs’ reasoning performance. Subsequent research has further refined the methods for eliciting rationales, leading to improved performance (Fu et al., 2023; Zhou et al., 2022).
Figure 1: A simplified example showing how implicit rationales in pre-training data can be leveraged to improve reasoning. 1: Implicit rationales (unstated logical connections) occur frequently in LLM pre-training data. 2: As a result, existing LLMs pre-trained to replicate their pretraining data tend to omit these logical steps as well. 3: However, Rationalyst learns to generate these rationales at inference time to supervise the chain-of-thought process for more accurate reasoning.
In the context of LLM reasoning, these rationales are typically employed through a chain-of-thought process that makes reasoning steps explicit by articulating them as plain-text rationales. In this approach, each subsequent rationale is generated based on rationales produced in preceding steps, effectively using them as a form of supervision. However, the generated reasoning chains might be incomplete, containing potential logical leaps while leaving some rationales implicit (or hidden) during the generation process. These gaps in the reasoning chain can weaken the LLM’s reasoning ability throughout the problem-solving process.
@@
This paper presents Rationalyst, a model tailored for process-supervision of reasoning. Rationalyst is pre-trained on a vast collection of implicit rationales extracted from a mixture of web-scale unlabeled datasets and existing reasoning datasets. Although existing LLMs may miss crucial details in their reasoning, leading to flawed conclusions ( 2 in Figure 1), integrating these LLMs with Rationalyst provides an additional supervision mechanism to guide their reasoning processes, resulting in more robust conclusions ( 3 in Figure 1).
@@
Our method extracts 65k implicit rationales from the web-scale unlabelled dataset The Pile (Gao et al., 2020). To adapt the extracted rationales to our tested domain and stabilize training, we additionally extract a much smaller set of 14k implicit rationales from the question-answer pairs in the training sets of two reasoning datasets: GSM8K (Cobbe et al., 2021a) and ECQA (Aggarwal et al., 2021). Our extraction process controls for answer leakage to prevent artificial amplification of performance. Using this curated set of rationales, Rationalyst is then fine-tuned from LLaMa-3-8B. To assess the effectiveness of our approach, we evaluate Rationalyst on a diverse set of reasoning tasks, including mathematical, commonsense, scientific, and logical reasoning. Our results show that Rationalyst improves the accuracy of reasoning by an average of 3.9% (§5.1). To understand the contribution of different data sources, we conduct an ablation study that demonstrates the utility of rationales from both the large-scale Pile dataset and the smaller, specialized reasoning datasets (§5.2). Notably, Rationalyst exhibits superior performance when compared to strong general-purpose verifiers like GPT-4 and similar capacity models specifically fine-tuned on matching training sets (§5.4).
Implicit rationales generated by Rationalyst are also designed to provide supervision in a human-readable form, offering improved interpretability for LLM generation. This added interpretability is particularly beneficial when reasoning over complex domains such as mathematics or coding, where the step-by-step logic can be difficult for humans to follow without explicit explanations. As shown in §5.5, our model is capable of generating human-understandable rationales for unseen data from complex math reasoning.
Our contributions in this paper are two-fold:
  • We propose Rationalyst, a model that is pre-trained on implicit rationales extracted from unlabeled text data. Rationalyst enhances LLM interpretability and performance during reasoning by providing process supervision.
  • We empirically show Rationalyst generalizes across reasoning tasks and scales with unlabelled data.
@@ 2 Related Work
Supervision-based approaches have been shown to enhance the reasoning abilities of LLMs. Cobbe et al. (2021b) and Snell et al. (2024) demonstrate that training a “verifier" to supervise reasoning can be more parameter-efficient than simply expanding the parameters of the “reasoner" responsible for solving the reasoning task. Ground-truth feedback from interaction with the environment is an effective form of supervision (Wang et al., 2023), but it works only in controlled environments like simulated world. General-purpose verifiers (Dhuliawala et al., 2023; Weir et al., 2024, 2023; Vacareanu et al., 2024) offer broader applicability utilizing principles like compositional reasoning. However, they don’t fully capitalize on the vast amount of unlabelled data in the way a data-driven approach might. Process-based supervision (Lightman et al., 2023) offers supervision at each reasoning step rather than just at the final result. While promising, it requires substantial human annotation for the correctness of intermediate steps, making it resource-intensive. Our work aims to address these challenges by proposing a data-centric process-supervision method without the need for human annotation.
@@ 3.1 Large-scale Rationale Extraction
Implicit rationales are often embedded in unlabelled text, reflecting natural thought processes in daily communication. Our extraction process, illustrated in Figure 3, aims to make these rationales explicit. Using a pre-trained and aligned language model $M$, we generate rationales from text and then use $M$ to filter these rationales to retain only those that are useful, akin to the self-supervised “tool” learning approach described by Schick et al. (2023). The same $M$ is subsequently used to train Rationalyst.
@@
Extracting rationales from pre-training data. We employ $M$ to generate rationales from the Pile. Due to the size of this dataset, we implement a pre-filtering process to identify reasoning-rich documents by (1) computing the average semantic embedding of representative reasoning training sets using a paragraph embedding model, and (2) selecting documents from unlabelled datasets that exceed a cosine similarity threshold $\alpha$ when compared to this average embedding. After pre-filtering, we segment the selected paragraphs into 2000-word segments and instruct $M$ to generate rationales at the end of each sentence, using prompts with demonstrations. Detailed information on the prompts and in-context learning demonstrations used for rationale extraction can be found in Appendix A.
@@
Filtering extracted rationales. Generated rationales in A and B may not always be accurate or helpful. In reasoning tasks, our objective is for the extracted rationales to effectively aid in future reasoning, which means a good rationale should enhance the likelihood of accurately predicting the following text. Let $i$ be the position of the rationale $r$ in the sequence $\mathbf{x}=x_{1},\ldots,x_{n}$. Given a sequence of weights $(w_{k})_{k\in\mathbb{N}}$, the weighted cross-entropy loss for future token prediction is defined as:
@@
It’s crucial to clarify two key aspects of our rationale extraction process. First, while $M$ extracts rationales from the training sets of reasoning datasets, these training sets are not directly used as targets when training $M$ itself. Second, we explicitly instruct $M$ to exclude answers from the extracted rationales. This precaution prevents answer leakage in our prompts, thereby ensuring the integrity of our reasoning process.
@@ 3.2 Rationalyst Training
Given the appropriate context from either source, the implicit rationales, extracted and filtered as described in §3.1, serve as the target outputs during training. The overall training objective is to minimize the per-token cross-entropy loss between the generated rationales and their ground truth values from the extracted and filtered rationales. By learning to generate appropriate rationales for both free-form text and structured problem-solving data, Rationalyst develops the ability to provide meaningful guidance across a wide range of contexts during inference.
@@ 3.3 Inference with the Help of Rationalyst
During inference, a general-purpose LLM (the “agent model" or $M_{\text{Agent}}$) is employed for reasoning across various problems. Algorithm 1 outlines the procedure.
@@
For this supervision, $M_{\text{Agent}}$ generates the next reasoning steps conditioned on the trajectory $T$ (line 6). We then use $M$ to estimate the probability of potential next reasoning steps given rationale $r$ and reasoning trajectory $T$ (line 13). This probability-based heuristic aligns with our rationale filtration process used during $M_{\text{Ra}}$ training. Just as we identified rationales that improved the prediction of future text during filtration, here we use rationales to improve the selection of future reasoning steps. By leveraging the probability estimates as a heuristic, we can effectively discriminate between more and less likely next steps in the reasoning process, guiding the overall trajectory towards more accurate conclusions.
@@
6:     $r\leftarrow{\color[rgb]{0.9803,0.7372,0.0352}\definecolor[named]{pgfstrokecolor}{rgb}{0.9803,0.7372,0.0352}M_{\text{Ra}}}(T)$triangleright Generate implicit rationale given trajectory.
7:     $heuristic\_list=\emptyset$$\triangleright$ Empty the heuristic list at every step.
@@
9:         $next\_steps\leftarrow{\color[rgb]{0.3333,0.3764,0.6627}\definecolor[named]{pgfstrokecolor}{rgb}{0.3333,0.3764,0.6627}M_{\text{Agent}}}(T)$triangleright Sample next reasoning steps.
@@
16:              $h\leftarrow{\color[rgb]{0.9450,0.1,0.1}\definecolor[named]{pgfstrokecolor}{rgb}{0.9450,0.1,0.1}M}(x|T,r)$triangleright Estimate prob. of next reasoning step.
@@
20:         $heuristic\_list.append(h)$trianglerightRetain the heuristic.
@@
22:     $max\_idx\leftarrow argmax(heuristic\_list)$
23:     $T\leftarrow T+next\_steps[max\_idx]$triangleright Extend trajectory with the highest scoring step.
@@ 4.1 Setup for Training Rationalyst
Following the recipe in §3.1 B, we also extract rationales from existing reasoning datasets. GSM8K (Cobbe et al., 2021b) and ECQA (Aggarwal et al., 2021) were selected for their complementary coverage of mathematical and commonsense reasoning, respectively. This combination ensures Rationalyst is trained on diverse reasoning patterns, enhancing its versatility across various tasks.
@@
The model $M$ used for rationale extraction and rationale filtering are both LLama-3-8B-Instruct MetaAI (2024). On GSM8K and ECQA, we manually annotated 100 pairs of {preceding_context, rationale, following_context} to determine an appropriate filtration threshold. The annotations include 50 positive and 50 negative rationale examples. Since it’s straightforward to scale up the extraction of rationales from unlabelled data for filtration, we prioritize maximizing the precision of our filtered rationales, even if it means extracting fewer of them. We set the threshold $\tau_{f}$ to ensure that 95% of the filtered rationales are accurate. On The Pile, we do not perform rationale annotation due to its diverse composition of corpora with varying characteristics. So the filter threshold $\tau_{f}$ for the Pile is set to 0 for all of its subdomains.
@@
The resulting data from extraction/filteration.
The results of rationale extraction and filtration on GSM8K, ECQA, and The Pile are presented in Table 1. On GSM8K, our method generates an average of 2.34 rationales per document, while on ECQA, it generates 2.58 rationales per document. The filtration process removes 80.5% of the generated rationales on GSM8K and 42.4% on ECQA.
@@
Rationalyst is fine-tuned with LLaMa-3-8B-Instruct as the base model. We use the default hyperparameters as specified in the LLaMa-3 technical report (MetaAI, 2024) for fine-tuning.

### v2

# for Process Supervision of Reasoning

@@ Abstract
The reasoning steps generated by LLMs might be incomplete, as they mimic logical leaps common in everyday communication found in their pre-training data: underlying rationales are frequently left implicit (unstated). To address this challenge, we introduce Rationalyst, a three-stage approach for improving LLM reasoning: (1) extracting implicit rationales from unlabeled data, (2) training a specialized rationale generation model, and (3) using these rationales to provide process supervision during inference. We extract 79k rationales from web-scale unlabelled dataset (the Pile) and a combination of reasoning datasets with minimal human intervention. This web-scale pre-training for reasoning allows Rationalyst to consistently generalize across diverse reasoning tasks, including mathematical, commonsense, scientific, and logical reasoning. Fine-tuned from LLaMa-3-8B-Instruct, Rationalyst improves the accuracy of reasoning by an average of 3.9% on 7 representative reasoning benchmarks. It also demonstrates superior performance compared to significantly larger verifiers like GPT-4 and similarly sized models fine-tuned on matching training sets. 111 Our code, data, and model can be found at this repository: https://github.com/JHU-CLSP/RATIONALYST
ACL’25
[Uncaptioned image] Rationalyst: Mining Implicit Rationales
for Process Supervision of Reasoning
@@ 1 Introduction
Rationales play a crucial role in human reasoning and its accuracy (Rips, 1994; Mercier and Sperber, 2011). In reasoning problems, having accurate rationales often correlates with accurate outcomes (Tversky et al., 1982; Davis, 1984). This importance of rationales extends to Large Language Models (LLMs) as well. Wei et al. (2022) were among the first to show that generating chain-of-thought rationales significantly improves LLMs’ reasoning performance. Subsequent research has further refined the methods for eliciting rationales, leading to improved performance (Zhou et al., 2022).
Figure 1: A simplified example showing how implicit rationales in pre-training data can be leveraged to improve reasoning. 1: Implicit rationales (unstated logical connections) occur frequently in LLM pre-training data. 2: Existing LLMs pre-trained to replicate their pretraining data tend to omit these logical steps as well. 3: However, Rationalyst learns to generate these rationales at inference time to supervise the chain-of-thought process for better reasoning.
In the context of LLM reasoning, these rationales are typically employed through a chain-of-thought process that makes reasoning explicit by articulating them as plain-text rationales. In this approach, each subsequent rationale is generated conditioned on rationales produced in preceding steps, effectively using them as a form of supervision. However, the generated reasoning chains might be incomplete, containing potential logical leaps while leaving some rationales implicit (or hidden) during the generation process. These gaps in the reasoning chain can weaken the LLM’s reasoning ability throughout the problem-solving process.
@@
This paper presents Rationalyst, a model tailored for process-supervision of reasoning. Rationalyst is trained on a vast collection of implicit rationales extracted from a mixture of web-scale unlabeled datasets and existing reasoning datasets. Although existing LLMs may miss crucial details in their reasoning, leading to flawed conclusions ( 2 in Figure 1), using Rationalyst provides additional supervision mechanism to guide their reasoning processes, resulting in more robust conclusions ( 3 in Figure 1).
@@
Our method extracts 65k implicit rationales from the web-scale unlabelled dataset The Pile (Gao et al., 2020). To adapt the extracted rationales to our tested domain and stabilize training, we additionally extract a much smaller set of 14k implicit rationales from the question-answer pairs in the training sets of two reasoning datasets: GSM8K and ECQA. Our extraction process controls for answer leakage to prevent artificial amplification of performance. Using this curated set of rationales, Rationalyst is then fine-tuned from LLaMa-3-8B. To assess the effectiveness of our approach, we evaluate Rationalyst on a diverse set of reasoning tasks, including mathematical, commonsense, scientific, and logical reasoning. Our results show that Rationalyst improves the accuracy of reasoning by an average of 3.9% (§5.1). To understand the contribution of different data sources, we conduct an ablation study that demonstrates the utility of rationales from both the large-scale Pile dataset and the smaller, specialized reasoning datasets (§5.2). Notably, Rationalyst exhibits superior performance when compared to strong general-purpose verifiers like GPT-4 and similar capacity models specifically fine-tuned on matching training sets (§5.4).
Implicit rationales generated by Rationalyst are also designed to provide supervision in a human-readable form, offering improved interpretability for LLM generation. This added interpretability is particularly beneficial when reasoning over complex domains such as mathematics or coding, where the step-by-step logic can be difficult for humans to follow without explicit explanations. As shown in §5.5, our model is capable of generating human-understandable rationales for unseen data from complex math reasoning. Furthermore, our analysis demonstrates that Rationalyst introduces minimal computational overhead during inference (§5.6).
@@ 2 Related Work
Supervision-based approaches have been shown to enhance the reasoning abilities of LLMs. Cobbe et al. (2021) and Snell et al. (2024) demonstrate that training a “verifier" to supervise reasoning can be more parameter-efficient than simply expanding the parameters of the “reasoner" responsible for solving the reasoning task. Ground-truth feedback from interaction with the environment is an effective form of supervision (Wang et al., 2023), but it works only in controlled environments like simulated world. General-purpose verifiers (Dhuliawala et al., 2023; Weir et al., 2024, 2023; Vacareanu et al., 2024) offer broader applicability utilizing principles like compositional reasoning. However, they don’t fully capitalize on the vast amount of unlabelled data in the way a data-driven approach might. Process-based supervision (Lightman et al., 2023) offers supervision at each reasoning step rather than just at the final result. While promising, it requires substantial human annotation for the correctness of intermediate steps, making it resource-intensive. Our work aims to address these challenges by proposing a data-centric process-supervision method without the need for human annotation.
@@
LLMs are conventionally trained on extensive web data using autoregressive next-token prediction. While effective, this approach may not fully harness the potential of the pre-training data, as latent information within this data could be better accessed using techniques beyond simple next-token prediction. Recent research has demonstrated several approaches to utilize this latent information to develop more sophisticated language model capabilities (Jiang et al., 2024c). Schick et al. (2023) introduced Toolformer, which autonomously annotates and extracts appropriate positions, names, and inputs for tool use by leveraging supervision from future tokens. Similarly, Cornille et al. (2024) developed a method for learning to plan coherent article writing through self-supervised learning in text. More closely related to our work, Zelikman et al. (2024) proposed Quiet-Star, which applied a comparable technique to uncover underlying rationales in daily communication to enhance reasoning capabilities. Our work adopts a strategy similar to Quiet-Star for extracting rationales in an unsupervised manner. However, our approach diverges in its primary objective: we aim to train a “supervisor" that can utilize these rationales to provide process supervision for any “reasoner." This focus enables us to implement a simpler and more reliable method, as we don’t need to directly integrate rationale extraction with “reasoner" training. Our approach thus offers a novel perspective on leveraging latent information in language models to enhance their capabilities.
@@ 3.1 Large-scale Rationale Extraction
Implicit rationales are often embedded in unlabelled text, reflecting natural thought processes in daily communication. Our extraction process, illustrated in Figure 3, aims to make these rationales explicit. Using an aligned language model $M$, we generate rationales from text and then use $M$ to filter these rationales to retain only those that are useful, akin to the self-supervised “tool” learning approach described by Schick et al. (2023). The same $M$ is subsequently used to train $M_{\text{Ra}}$.
@@
Extracting rationales from pre-training data. We employ $M$ to generate rationales from the Pile. Due to the size of this dataset, we implement a pre-filtering process to identify reasoning-rich documents by (1) computing the average semantic embedding of representative reasoning training sets using a paragraph embedding model, and (2) selecting documents from unlabelled datasets that exceed a cosine similarity threshold $\alpha$ when compared to this average embedding. After pre-filtering, we segment the selected paragraphs into 2000-word segments and instruct $M$ to generate rationales at the end of each sentence, using prompts with demonstrations. Detailed information on the prompts and in-context learning demonstrations used for rationale extraction can be found in Appendix B.
@@
Filtering extracted rationales. Generated rationales in A and B may not always be accurate or helpful. In reasoning tasks, our objective is for the extracted rationales to effectively aid in future reasoning, which means a good rationale should enhance the likelihood of accurately predicting the following text. Let $i$ be the position of the rationale $r$ in the sequence $\mathbf{x}=x_{1},\ldots,x_{n}$. Given a sequence of weights $(w_{k})_{k\in\mathbb{N}}$, the cross-entropy loss for future token prediction is defined as:
@@
It’s crucial to clarify two key aspects of our rationale extraction process. First, while $M$ extracts rationales from the training sets of reasoning datasets, these training sets are not directly used as targets when training $M_{\text{Ra}}$. Instead, the target of training is the rationale generated by $M$. Second, we explicitly instruct $M$ to exclude answers from the extracted rationales. This precaution prevents answer leakage in our prompts.
@@ 3.2 Rationalyst Training
Given the appropriate context from either source, the implicit rationales, extracted and filtered as described in §3.1, serve as the target outputs during training. The overall training objective is to minimize the per-token cross-entropy loss between the generated rationales and their ground truth values from the extracted and filtered rationales.
@@ 3.3 Inference with the Help of Rationalyst
During inference, any general-purpose LLM (the “agent model" or $M_{\text{Agent}}$) can be employed for reasoning across various problems. Algorithm 1 outlines the procedure.
@@
For this supervision, $M_{\text{Agent}}$ generates the next reasoning steps conditioned on the trajectory $T$ (line 6). We then use $M$ to estimate the probability of potential next reasoning steps given rationale $r$ and reasoning trajectory $T$ (line 13). This probability-based heuristic aligns with our rationale filtration process used during $M_{\text{Ra}}$ training: just as we identified rationales that improved the prediction of future text during filtration, here we use rationales to improve the selection of future reasoning steps. By leveraging the probability estimates as a heuristic, we can discriminate between more and less likely next steps in the reasoning process, guiding the overall trajectory towards more accurate conclusions.
@@
6:     $r\leftarrow{\color[rgb]{0.9803,0.7372,0.0352}\definecolor[named]{pgfstrokecolor}{rgb}{0.9803,0.7372,0.0352}M_{\text{Ra}}}(T)$triangleright Generate rationale given trajectory.
7:     $heuristic\_list=\emptyset$
@@
9:         $next\_steps\leftarrow{\color[rgb]{0.3333,0.3764,0.6627}\definecolor[named]{pgfstrokecolor}{rgb}{0.3333,0.3764,0.6627}M_{\text{Agent}}}(T)$
@@
16:              $h\leftarrow{\color[rgb]{0.9450,0.1,0.1}\definecolor[named]{pgfstrokecolor}{rgb}{0.9450,0.1,0.1}M}(x|T,r)$triangleright Estimate probability of next reasoning step.
@@
20:         $heuristic\_list.append(h)$
@@
22:     $m\_idx\leftarrow argmax(heuristic\_list)$
23:     $T\leftarrow T+next\_steps[m\_idx]$triangleright Extend trajectory with the highest scoring step.
@@ 4.1 Setup for Training Rationalyst
Following the recipe in §3.1 B, we also extract rationales from existing reasoning datasets. GSM8K (Cobbe et al., 2021) and ECQA (Aggarwal et al., 2021) were selected for their complementary coverage of mathematical and commonsense reasoning, respectively. This combination ensures Rationalyst is trained on diverse reasoning patterns, enhancing its versatility across various tasks.
@@
The model $M$ used for rationale extraction and rationale filtering are both LLama-3-8B-Instruct MetaAI (2024). On GSM8K and ECQA, we manually annotated 100 pairs of {preceding_context, rationale, following_context} to determine an appropriate filtration threshold. The annotations include 50 positive and 50 negative rationale examples. Since it’s straightforward to scale up the extraction of rationales from unlabelled data for filtration, we prioritize maximizing the precision of our filtered rationales, even if it means extracting fewer of them. We set the threshold $\tau_{f}$ to ensure that 95% of the filtered rationales are accurate. On The Pile, we do not perform rationale annotation due to its diverse composition of corpora with varying characteristics. So the filter threshold $\tau_{f}$ for the Pile is set to 0 for all of its subdomains. The results of rationale extraction and filtration on GSM8K, ECQA, and The Pile are presented in Table 1. From our extraction and filtration process, we obtained approximately 14k rationales from GSM8K and ECQA combined, and about 65k from The Pile after filtration. Detailed statistics of rationale extraction and filtration results for each dataset, including the number of rationales per document and filtration rates, can be found in Appendix G.
@@
Rationalyst is fine-tuned with LLaMa-3-8B-Instruct as the base model. We use the default hyperparameters as specified in the LLaMa-3 technical report (MetaAI, 2024) for fine-tuning. After training, we conducted manual annotation of the model’s output and found that the accuracy of rationales generated on unseen test data closely matches the filtration accuracy we specified for training data through our filtration parameters.